OpenStudy (anonymous):

Dinitrogen pentoxide can be made by the reaction of nitrogen with oxygen. Consider the reaction of 35g of nitrogen with 48 grams of oxygen.What is the maximum number of grams of dinitrogen pentoxide that can be made?

OpenStudy (anonymous):

I got 270 grams as my answer but the answer is 65 grams. I'm confused

OpenStudy (anonymous):

2N2+5O2 --> 2(N2O5) balance the equation first convert 35g N2 to moles convert 48g O2 to moles find the limiting reagent (which ever has less moles), then mole-ratio to 2(N2O5) and find your mass, hope this helped!

OpenStudy (anonymous):

I converted it and got 2.5 mole N2 and 3 moles O2. I found teh limiting reagent to ve N2 since it produces 2.5 moles N205. Then i multiplied 2.5 moles * 108 (the molecular weight) and got 270. Im not sure what I'm doing wrong.

OpenStudy (anonymous):

Strange, I got 135g of N2O5

OpenStudy (anonymous):

Wow all three of the answers are very different. The answer key says 65, I got 270 and you got 135

OpenStudy (anonymous):

make sure you're using N2 molar mass of (14.01*2) and O2 molar mass of (16*2) 35g/28.02g/mol = 1.25 moles N2 48g/32g/mol = 1.406 moles O2 N2 = limiting reagent mole ratio 1.25 mole N2 = 1.25 mole N2O5 = 1.25 mol * molar mass of N2O5 (14.01*2+16*5) = 135.025 Im pretty sure I did the math right

OpenStudy (anonymous):

Ok im going to try it myself also becuase i forgot N2 AND O2 were diatomic. Ughh but why the asnwer is 65...idk :(

OpenStudy (anonymous):

Sometimes answer key could be wrong, I really don't see where I could've miscalculated, I've triple checked my work lol

OpenStudy (anonymous):

lol thanks :) and yeah i got 135 g too

OpenStudy (anonymous):

i think 65 is incorrect. thanks. I'll ask my professor if tehre was a glitch in the asnwer key.

OpenStudy (anonymous):

but the weird thing is that when you multiply 108 g by 0.6 mol O2 you actually get 64.8 or 65 grams of N2O5. Are we both not seeing how O2 could possibly be the limiting reagent

OpenStudy (anonymous):

But the moles of O2 was more than N2 which made the N2 the limiting reagent

OpenStudy (anonymous):

Thats what I thought too...till someone on a discussuion borad wrote O2 is the limiting reagent b/c the ratio of O2 to N2O5 IS 2/5 and the ratio of N2 to N2O5 is 2/2. Does taht make any sense to you?

OpenStudy (anonymous):

Ah yes, im sorry, O2 should be the limiting reagent :), forgot to divide because the molar ratios were not 1 to 1. Forgive me it's almost midnight here :)

OpenStudy (anonymous):

Yes its almost 1am here too! And sorry to bother you but can you please explain why in this situation we dont look at the moles we ended up with to find the LR but teh ratio of the reactants to teh products?

OpenStudy (anonymous):

Limiting reagent you have to divide the moles to find the true ratio

OpenStudy (anonymous):

idk what you mean?? srry :/